# Seeding with Differentially Private Network Information

### v2

@@
When designing interventions in public health, development, and education, decision makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population. Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for designing sustainable public health and development interventions on social networks. In a similar vein, social media platforms rely on network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising. Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies. We study privacy guarantees for influence maximization algorithms when the social network is unknown, and the inputs are samples of prior influence cascades that are collected at random. Building on recent results that address seeding with costly network information, our privacy-preserving algorithms introduce randomization in the collected data or the algorithm output, and can bound each node’s (or group of nodes’) privacy loss in deciding whether or not their data should be included in the algorithm input. We provide theoretical guarantees of the seeding performance with a limited sample size subject to differential privacy budgets in both central and local privacy regimes. Simulations on synthetic and empirical network datasets reveal the diminishing value of network information with decreasing privacy budget in both regimes.
@@
social network interventions, differential privacy, influence maximization, sample complexity, approximation algorithms, peer influence and social contagion, public health interventions
@@ 1 Introduction
When designing interventions in public health, development, and education decision-makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population. Many of these seeding methods are implemented without complete knowledge of social network structures — e.g., by targeting nominated friends of random individuals (Kim et al. 2015) or random pairs of friends (Alexander et al. 2022) — and they enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group). However, even without complete network information, network participants (and in particular minority members) who contribute to partial social network data collection endure privacy risks. This is especially important when engaging with vulnerable populations to record sensitive or stigmatizing behavior, e.g., in design of public health interventions among the homeless youth for HIV or suicide prevention (Wilder et al. 2021). Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for responsible design of public health interventions on social networks and maintaining community engagements that sustain such interventions.
@@
In this paper we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability. Under these schemes the network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations. We investigate how the targeting efficacy is affected by decreasing the privacy budget that entails more noise in the collected data (with local privacy mechanisms) or the targeting algorithm’s output (with central privacy mechanisms).
@@
After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to social network data privacy for seeding interventions and motivate our proposed notion of “influence sample differential privacy”. In Section 3, we provide privacy mechanisms for seeding in both central (Section 3.1) and local (Section 3.2) regimes. In Section 4, we discuss and contextualize the nature of our privacy protections for social network data collection and intervention design applications. In Section 5, we test the performance of our privacy mechanisms on synthetic and real social network datasets. We end with a discussion of limitations, future directions and concluding remarks in Section 6. Complete proofs and supplementary results are provided in several appendices at the end.
@@ 1.1 Summary of the main results
Differential privacy (DP) offers an information-theoretic foundation for privacy by focusing on protecting input data against revelations of algorithmic outputs (Dwork et al. 2014). In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in its input domain and any subset $A$ in its output range we have:
@@
The $\epsilon$-DP criterion in Equation 1 limits inferences of any observer/adversary about the input data based on the algorithm output. The adjacency notion between datasets is typically defined such that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference. Hence, different notions of adjacency will imply different protections for the individuals. In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”), and the nature of privacy protections that it offers. Throughout this paper we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without hyphen between the words as appropriate by the language function that these words play in the context that DP is used.
@@
We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum under a randomized model of network diffusion, whereby each edge passes on the influence (or “infection”) with an independent cascade probability $p$. Given the social network graph ${\mathcal{G}}$ and an integer $k\geq 1$, we denote the expected spread size from infecting the nodes in $S$ by $I_{\mathcal{G}}(S)$ and write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{\mathcal{G}}(S)$ for the optimum solution to the $k$-influence maximization problem. It is an example of a cardinality-constrained combinatorial optimization on a graph with a monotone, submodular objective that admits a tight $(1-1/e)$ approximation guarantee using the greedy node selection algorithm. This problem has occupied a central place in graph and social networks algorithms literature since its introduction by Kempe et al. (2003). There is a wealth of literature (spanning almost two decades) on efficient approximations of this problem (Kempe et al. 2003, 2005, 2015, Chen et al. 2009, Wang et al. 2012, Borgs et al. 2014). However, the graph structure of the input data has stunned the design of DP influence maximization algorithms. Common operationalizations of DP for graph algorithms, such as node- or edge-DP define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge. This adjacency criteria is often too stringent for graph problems whose solutions are extremely sensitive to existence of nodes and edges (see Section 1.2 for detailed citations and comparisons). Our main hardness results indicate that it is impossible to obtain useful performance guarantees for influence maximization subject to node- or edge-DP constraints (see Section 2.2, LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg, summarized in 1 below).
@@
Given $n\geq 1$ and $\epsilon>0$, there can be no $\epsilon$-node-DP or $\epsilon$-edge-DP seeding algorithms that give an approximation guarantee better than $(1-1/e)\mbox{\sf OPT}-\alpha_{\epsilon}{n}$ for $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$.
@@
Our main contribution here is to formulate a new differential privacy notion with guarantees for influence maximization when the social network is unknown and the algorithm inputs are samples of prior influence cascades that are collected at random, i.e., “influence samples” — rather than the entire social network graph. Our results build on recent work about seeding with costly network information (Eckles et al. 2022), and more broadly sample complexity of influence maximization (Sadeh et al. 2020, Borgs et al. 2014). Our algorithm input is a set of $m$ influence samples, i.e., $m$ independent cascade samples that are collected on the social network graph each starting from a random, initially-infected node. We denote this input deadset by a binary matrix ${\mathbf{x}}\in\{0,1\}^{n\times m}$, where each row represents a node and each column represents the status of the node in an influence sample: $[{\mathbf{x}}]_{ij}=1$ if node $i$ is infected in the $j$-th influence sample and $[{\mathbf{x}}]_{ij}=0$ otherwise. Our privacy notion is with respect to the contribution of individual network participants to the influence samples, hence, we define the adjacent input datasets in Equation 1 as ${\mathbf{x}},{\mathbf{x}}^{\prime}\in\{0,1\}^{n\times m}$ which differ in a single binary entry:
@@
We follow the formulation of influence maximization by Eckles et al. (2022) where the social network is unknown and influential nodes need to be estimated from costly samples of past spreads. Our privacy guarantees are with respect to influence samples, controlling information leakage about the cascade data. For example, Apple Inc. uses a randomized response mechanism to collect the users’ daily usage of emojis with a fixed privacy budget ($\epsilon=4$). This setup maps naturally to our formulation of ISDP as the usage of each emoji can be modeled as an influence sample. We make no assumptions about adversary’s knowledge of the network and focus on differential privacy guarantees for influence maximization according to the ISDP formulation above.
@@
By introducing randomization at the input or output of seeding algorithms, we are able to bound each node’s privacy loss in deciding whether or not their data should be included in the influence samples — as opposed to the social network graph. Of note, this is critically different from DP analysis of graphs that is defined with respect to neighboring graphs in the algorithm input. In our approach, each node’s decision to appear in the input does not affect the data generation process (DGP) of the input sample or the underlying diffusion process; the network graph remains the same and the privacy implications are studied with respect to the construction of the samples on a fixed network. This allows us to meaningfully bound the effect of the removal of a node’s data on the algorithmic performance and to use those bounds to tradeoff the privacy risks against improved performance. To mitigate privacy risks, DP uses randomization to provide plausible deniability certificates to those contributing to the input data. This is achieved in one of the two ways: (i) using exponential mechanism (McSherry and Talwar 2007) by randomizing the algorithm output, i.e., central DP, (ii) using the randomized response mechanism (Warner 1965) by injecting noise in the input data, i.e., local DP. Accordingly, we propose two efficient DP seeding algorithms and bound their performance in terms of the sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$, (See Section 3 LABEL:thm:exp:acc and LABEL:thm:rr:acc):
@@
Given $m$ influence samples, in Section 3.1 we use exponential mechanisms (McSherry and Talwar 2007) to choose high quality seeds with central DP guarantees. In Subsection 3.2, before accessing the data (influence samples), we use the randomized response mechanism (Warner 1965), to perturb the input data and protect privacy at the local level. However, these perturbations can introduce additional biased noise into the data, so we provide a closed-form characterization of the introduced noise and an algorithm to offset this bias. In Section 5, we thoroughly evaluate our algorithms on synthetic and real graphs of different sizes and verify our theoretical guarantees in simulations.
@@ 1.2 Comparisons with related work
A long line of literature, going back to the seminal works of Kempe et al. (2003, 2005, 2015) propose efficient approximation algorithms for this problem based on the submodularity of the influence function (Chen et al. 2009, Wang et al. 2012). This problem arise naturally in applications such as development economics and public health to facilitate adoption of welfare-enhancing technologies, e.g., microfinance (Banerjee et al. 2013), insurance (Cai et al. 2015), contraception (Behrman et al. 2002), water purification methods, and multivitamins (Kim et al. 2015); and to target localized social network interventions for spreading information about immunization camps (Banerjee et al. 2019), preventing misinformation about drug side effects (Chami et al. 2017), increasing HIV awareness among homeless youth (Yadav et al. 2017), and reducing bullying among adolescent youth (Paluck et al. 2016). Alexander et al. (2022) show that network-based targeting substantially increases population-level adoption of iron-fortified salt to prevent iron-deficiency anemia among female household heads in India. Kim et al. (2015) show similar results for a cluster-randomized trial in villages in Honduras, where they measure the efficacy of social network targeting strategies to promote adoption of chlorine for water purification and multivitamins for micronutrient deficiencies. In all such settings, collecting complete social network connection (edge) data is difficult, costly, or impossible and comes with critical privacy and security risks to individuals in the social network. In this paper we focus on differential privacy guarantees for influence maximization when social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
@@
Much of the past work on social network data privacy focuses on identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (Romanini et al. 2021). Classical anonymization techniques mask node attributes and perturb, modify, randomize or aggregate the graph structure to prevent re-identification of nodes within a confidence level — typically achieving k-anonymity (Wu et al. 2010, Zheleva and Getoor 2011, Abawajy et al. 2016). However, statistical disclosure control of collected data requires safeguarding data donor’s privacy against adversarial attacks where common anonymization techniques are shown to be vulnerable to various kinds of identification (Barbaro et al. 2006), linkage and cross-referencing (Narayanan and Shmatikov 2008, Sweeney 2015, 1997), statistical difference and re-identification attacks (Kumar et al. 2007). On the other hand while edge sampling at the data collection can somewhat mitigate the re-identification risk of nodes (Romanini et al. 2021, Jiang et al. 2021), the seeded nodes can reveal additional side information about who has contributed to network data collection. For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks, even though no social network data are published. Our privacy guarantees for influence maximization follow the DP formalism to provide network participants and data donors with plausible deniability with regard to their contribution to the collected data or their role in shaping the algorithm output (the seeded nodes). DP is a gold standard for privacy-preserving data analysis and avoids the above drawbacks by abstracting away from the attack models and controlling the privacy leak at an information-theoretic level according to Equation 1.
@@
Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms, and it has been impactful in large-scale and high-profile applications such as the U.S. decennial census Garfinkel (2022) and the tech industry Erlingsson (2014a), Apple Differential Privacy Team (2017). It have been successfully applied to a broad variety of problem classes, including, e.g., combinatorial (Gupta et al. 2010), submodular (Mitrovic et al. 2017, Sadeghi and Fazel 2021), and online submodular (Salazar and Cummings 2021) optimization problems. The existing literature on DP analyses of graphs, e.g., edge-DP Nissim et al. (2007) or node-DP (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016), has devoted much attention to graph statistics and parameters such as average degree and various subgraph counts (Blocki et al. 2022), degree distribution (Hay et al. 2009) and the graph spectra (Wang et al. 2013), as well as community Mohamed et al. (2022) and densest subgraph detection (Nguyen and Vullikanti 2021). These techniques are not directly applicable to influence maximization, and our main hardness results indicate that node- or edge-DP are too stringent to render useful performance guarantees for influence maximization. Similar observations have been made about node- and edge-DP guarantees of other graph algorithms which are too sensitive to the presence of individual edges (and especially nodes) in their input graphs (Gupta et al. 2010, Li et al. 2022, Epasto et al. 2022, Blocki et al. 2013). In their analysis of deferentially private combinatorial optimization, Gupta et al. (2010) consider several graph problems including vertex and set cover. Their modified greedy algorithm for set cover is similar to our central DP mechanism in Section 3.1, though addressing a substantially different privacy environment, where the goal is to protect a private subset $R$ while outputting a minimal collection that covers $R$. In both cases direct application of the exponential mechanism is inefficient and the greedy approach facilitates an efficient implementation.
@@
In formulating our notion of “influence sample differential privacy” (ISDP), we move away from having the social network as the algorithm input. Instead, we focus on an input dataset of $m$ cascades that are collected at random on the social network graph. ISDP allows us to limit the information leakage about contribution of individual nodes to the input cascades (with local ISDP), or contribution of individual nodes to the algorithm output through the input cascades (with central ISDP). It is worth noting that achieving local privacy under ISDP is easy by randomizing individual statuses of the nodes in the input cascades. However, in general it is challenging to coordinate randomization across nodes in a meaningfully decentralized manner that matches the problem context, and locally deferentially private analysis of graphs has been limited so far to certain restricted settings (Imola et al. 2021, Dhulipala et al. 2022).
@@
The fundamental relationship between information flow and privacy is well noted, cf., contextual integrity (Nissenbaum 2009, Benthall and Cummings 2022). Notwithstanding, the privacy implications of information diffusion and algorithmic intervention on social networks is largely unexplored, with the exception of two studies (Rezaei et al. 2021, Liell-Cock et al. 2020). The existing work looks at this issue in specific, highly-stylized scenarios. One study shows that it is extremely hard to hide the existence of a giant, connected component of infected nodes under an independent cascade model and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al. 2021). In another study, the authors propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al. 2020). With our formulation of DP for influence maximization we can focus on privacy leaks through information flow on the network rather than the underlying network structure. Our formulation of DP for influence samples ($\epsilon$-ISDP in 1) allows us to explicitly quantify information leaks from observation of random spreads on the network and for us to take a direct approach to the privacy risks that social network participants face from network diffusions.

### v3

@@
In public health interventions such as the distribution of preexposure prophylaxis (PrEP) for HIV prevention, decision makers rely on seeding algorithms to identify key individuals who can amplify the impact of their interventions. In such cases, building a complete sexual activity network is often infeasible due to privacy concerns. Instead, contact tracing can provide influence samples, that is, sequences of sexual contacts without requiring complete network information. This presents two challenges: protecting individual privacy in contact data and adapting seeding algorithms to work effectively with incomplete network information. To solve these two problems, we study privacy guarantees for influence maximization algorithms when the social network is unknown and the inputs are samples of prior influence cascades that are collected at random and need privacy protection. Building on recent results that address seeding with costly network information, our privacy-preserving algorithms introduce randomization in the collected data or the algorithm output, and can bound the privacy loss of each node (or group of nodes) in deciding to include their data in the algorithm input. We provide theoretical guarantees of seeding performance with a limited sample size subject to differential privacy budgets in both central and local privacy regimes. Simulations on synthetic random graphs and empirically grounded sexual contacts of men who have sex with men reveal the diminishing value of network information with decreasing privacy budget in both regimes and graceful decrease in performance with decreasing privacy budget in the central regime. Achieving good performance with local privacy guarantees requires relatively higher privacy budgets that confirm our theoretical expectations.
@@
social networks, differential privacy, influence maximization, sample complexity, approximation algorithms, sexual behavior, HIV, public health
@@ 1 Introduction
When designing interventions in public health, development and education, decision makers rely on data from social networks to target a small number of people, leveraging peer effects and social contagion to deliver the greatest welfare benefits to the population. Many of these seeding methods are implemented without complete knowledge of social network structures – for example, by targeting nominated friends of random individuals (Kim et al. 2015) or random pairs of friends (Alexander et al. 2022) – and enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group). However, even without complete network information, network participants (and particularly minority members) who contribute to partial social network data collection face privacy risks. This is especially important when vulnerable populations are involved in recording sensitive or stigmatizing behavior, such as the design of public health interventions among homeless youth for HIV or suicide prevention (Wilder et al. 2021). In the context of HIV prevention, the distribution of pre-exposure prophylaxis (PrEP) can benefit from seeding algorithms that identify high-risk individuals based on data from social networks (Ross et al. 2016, Simpson and Gumel 2017). However, constructing a complete sexual activity network for this purpose is challenging due to the sensitivity of the sexual behavior data. Instead, contact tracing can provide influence samples, which capture sequences of sexual activity without requiring full network knowledge. This introduces two key challenges: protecting the privacy of individuals contributing to contract tracing data (i.e., influence samples) and adapting seeding algorithms to effectively leverage these partial data. Developing privacy-preserving methods for network data collection and targeted interventions is critical for responsible public health interventions in social networks, particularly in vulnerable communities such as homeless youth, who are estimated to have a six to twelve times higher risk of HIV than housed youth (Caccamo et al. 2017, Santa Maria et al. 2020). Such methods ensure that privacy concerns do not undermine the trust and community engagement necessary for successful public health initiatives.
@@
In this paper, we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability. Under these schemes, network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations. We investigate how the targeting efficacy is affected by decreasing the privacy budget that entails more noise in the collected data (with local privacy mechanisms) or the targeting algorithm’s output (with central privacy mechanisms).
@@
After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to the privacy of social network data for seeding interventions and motivate our proposed notion of “influence sample differential privacy”. In Section 3, we provide privacy mechanisms for seeding in central (Section 3.1) and local (Section 3.2) regimes. In Section 4, we discuss and contextualize the nature of our privacy protections for social network data collection and intervention design applications. In Section 5, we test the performance of our privacy mechanisms on synthetic Erdős-Rényi networks and empirically-grounded networks of men who have sex with men (MSM). We end with a discussion of limitations, future directions, and the concluding remarks in Section 6. Complete proofs and supplementary results are provided in several appendices at the end.
@@ 1.1 Summary of the main results
Differential privacy (DP) offers an information-theoretic foundation for privacy by focusing on protecting input data against revelations of algorithmic outputs (Dwork et al. 2014). In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in the input domain and any subset $A$ in the output range we have:
@@
The $\epsilon$-DP criterion in Equation 1 limits the inferences of any observer/adversary about the input data based on the algorithm output. The notion of adjacency between datasets is typically defined so that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference. Hence, different notions of adjacency will imply different protections for individuals. In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”) and the nature of privacy protections that it offers. Throughout this paper, we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without hyphen between the words as appropriate by the language function that these words play in the context that DP is used.
@@
We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum in a randomized network diffusion model, whereby each edge passes on the influence (or “infection”) with an independent cascade probability $p$. Given the social network graph ${\mathcal{G}}$ and an integer $k\geq 1$, we denote the expected spread size from infecting the nodes in $S$ by $I_{\mathcal{G}}(S)$ and write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{\mathcal{G}}(S)$ for the optimum solution to the $k$-influence maximization problem. It is an example of a cardinality-constrained combinatorial optimization on a graph with a monotone, submodular objective that admits a tight $(1-1/e)$ approximation guarantee using the greedy node selection algorithm. This problem has occupied a central place in the graph and social networks algorithms literature since its introduction by Kempe et al. (2003). There is a wealth of literature (spanning almost two decades) on efficient approximations of this problem (Kempe et al. 2003, 2005, 2015, Chen et al. 2009, Wang et al. 2012, Borgs et al. 2014). However, the graph structure of the input data has stunned the design of DP influence maximization algorithms. Common operationalizations of DP for graph algorithms, such as node or edge DP, define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge. This adjacency criterion is often too stringent for graph problems whose solutions are extremely sensitive to the existence of nodes and edges (see Section 1.2 for detailed citations and comparisons). Our main hardness results indicate that it is impossible to obtain useful performance guarantees for influence maximization subject to node- or edge-DP constraints (see Section 2.2, LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg, summarized in 1 below).
@@
Given $n\geq 1$ and $\epsilon>0$, there can be no $\epsilon$ node-DP or $\epsilon$ edge-DP seeding algorithms that give an approximation guarantee better than $(1-1/e)\mbox{\sf OPT}-\alpha_{\epsilon}{n}$ for $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$.
@@
Our main contribution here is to formulate a new differential privacy notion with guarantees for influence maximization when the social network is unknown and the algorithm inputs are samples of prior influence cascades that are collected at random, i.e., “influence samples” — rather than the entire social network graph. Our results are based on recent work on seeding with costly network information (Eckles et al. 2022), and more generally sample complexity of influence maximization (Sadeh et al. 2020, Borgs et al. 2014). Our algorithm input is a set of $m$ influence samples, i.e., $m$ independent cascade samples that are collected on the social network graph, each starting from a random, initially infected node. We denote this input data set by a binary matrix ${\mathbf{x}}\in\{0,1\}^{n\times m}$, where each row represents a node and each column represents the status of the node in an influence sample: $[{\mathbf{x}}]_{ij}=1$ if node $i$ is infected in the $j$-th influence sample and $[{\mathbf{x}}]_{ij}=0$ otherwise. Our notion of privacy is with respect to the contribution of individual network participants to the influence samples; hence, we define the adjacent input data sets in Equation 1 as ${\mathbf{x}},{\mathbf{x}}^{\prime}\in\{0,1\}^{n\times m}$, which differ in a single binary entry:
@@
We follow the formulation of influence maximization by Eckles et al. (2022) where the social network is unknown and the influential nodes must be estimated from costly samples of past spreads. Our privacy guarantees are with respect to influence samples, controlling information leakage about the cascade data. For example, Apple Inc. uses a randomized response mechanism to collect the users’ daily usage of emojis with a fixed privacy budget ($\epsilon=4$); see Apple Differential Privacy Team (2017). This setup maps naturally to our formulation of ISDP as the usage of each emoji can be modeled as an influence sample. We make no assumptions about adversary’s knowledge of the network and focus on differential privacy guarantees for influence maximization according to the ISDP formulation above.
@@
By introducing randomization at the input or output of seeding algorithms, we are able to bound each node’s privacy loss in deciding whether or not their data should be included in the influence samples, as opposed to the social network graph. Notably, this is critically different from DP analysis of graphs that is defined with respect to neighboring graphs in the algorithm input. In our approach, each node’s decision to appear in the input does not affect the data generation process (DGP) of the input sample or the underlying diffusion process; the network graph remains the same, and the privacy implications are studied with respect to the construction of the samples on a fixed network. This allows us to meaningfully bound the effect of the removal of a node’s data on the algorithmic performance and to use those bounds to tradeoff the privacy risks against improved performance. To mitigate privacy risks, DP uses randomization to provide plausible deniability certificates to those who contribute to the input data. This is achieved in one of the two ways: (i) using the exponential mechanism (McSherry and Talwar 2007) by randomizing the algorithm output, i.e., central DP; (ii) using the randomized response mechanism (Warner 1965) by injecting noise in the input data, i.e., local DP. Consequently, we propose two efficient DP seeding algorithms and bound their performance in terms of sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$, (see Section 3 LABEL:thm:exp:acc and LABEL:thm:rr:acc):
@@
Given $m$ influence samples, in Section 3.1 we use exponential mechanisms (McSherry and Talwar 2007) to choose high-quality seeds with central DP guarantees. In Section 3.2, before accessing the data (influence samples), we use the randomized response mechanism (Warner 1965), to perturb the input data and protect privacy at the local level. However, these perturbations can introduce additional biased noise into the data, so we provide a closed-form characterization of the introduced noise and an algorithm to offset this bias. In Section 5, we thoroughly evaluate our algorithms on real and synthetic graphs of different sizes and verify our theoretical guarantees in simulations.
@@ 1.2 Comparisons with related work
A long line of literature, going back to the seminal works of Kempe et al. (2003, 2005, 2015) proposes efficient approximation algorithms for this problem based on the submodularity of the influence function (Chen et al. 2009, Wang et al. 2012). This problem arises naturally in applications such as development economics and public health to facilitate the adoption of welfare-enhancing technologies, e.g., microfinance (Banerjee et al. 2013), insurance (Cai et al. 2015), contraception (Behrman et al. 2002), water purification methods and multivitamins (Kim et al. 2015); and to target localized social network interventions to spread information about immunization camps (Banerjee et al. 2019), prevent misinformation about drug side effects (Chami et al. 2017), increase awareness of HIV among homeless youth (Yadav et al. 2017, 2018), and reduce bullying among adolescents (Paluck et al. 2016). Alexander et al. (2022) show that network-based targeting substantially increases population-level adoption of iron-fortified salt to prevent iron-deficiency anemia among female household heads in India. Kim et al. (2015) show similar results for a cluster-randomized trial in villages in Honduras, where they measure the effectiveness of social network targeting strategies to promote the adoption of chlorine for water purification and multivitamins for micronutrient deficiencies. In all such settings, collecting complete social network connection (edge) data is difficult, costly, or impossible and comes with critical privacy and security risks to individuals in the social network. In this paper, we focus on differential privacy guarantees for influence maximization when social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
@@
Much of the previous work on social network data privacy focuses on the identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (Romanini et al. 2021). Classical anonymization techniques mask the attributes of the nodes and perturb, modify, randomize, or aggregate the structure of the graph to prevent reidentification of the nodes within a confidence level, typically achieving $k$ anonymity (Wu et al. 2010, Zheleva and Getoor 2011, Abawajy et al. 2016). However, statistical disclosure control of collected data requires safeguarding data donor’s privacy against adversarial attacks where common anonymization techniques are shown to be vulnerable to various kinds of identification (Barbaro et al. 2006), linkage and cross-referencing (Narayanan and Shmatikov 2008, Sweeney 2015, 1997), statistical difference and reidentification attacks (Kumar et al. 2007). However, while edge sampling at data collection can somewhat mitigate the risk of reidentification of nodes (Romanini et al. 2021, Jiang et al. 2021), seeded nodes can reveal additional side information about who has contributed to data collection. For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks even though no social network data is published. Our privacy guarantees for influence maximization follow the DP formalism to provide network participants and data donors with plausible deniability with regard to their contribution to the collected data or their role in shaping the algorithm output (the seeded nodes). DP is a gold standard for privacy-preserving data analysis and avoids the above drawbacks by abstracting away from the attack models and controlling the privacy leak at an information-theoretic level according to Equation 1.
@@
Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms and has been influential in large-scale and high-profile applications such as the US Census Garfinkel (2022) and the tech industry Erlingsson (2014a), Apple Differential Privacy Team (2017). It has been successfully applied to a broad variety of problem classes, including, e.g., shortest path (Fioretto et al. 2024), combinatorial (Gupta et al. 2010), submodular (Mitrovic et al. 2017, Sadeghi and Fazel 2021), and online submodular (Salazar and Cummings 2021) optimization problems. The existing literature on graph DP analysis, for example, edge DP (Nissim et al. 2007, Eden et al. 2023) or node DP (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016, Kalemaj et al. 2023), has devoted much attention to graph statistics and parameters such as average degree and various subgraph counts (Blocki et al. 2022), degree distribution (Hay et al. 2009) and graph spectra (Wang et al. 2013), as well as community detection Mohamed et al. (2022) and densest subgraph detection problems (Nguyen and Vullikanti 2021). These techniques are not directly applicable to influence maximization, and our main hardness results indicate that node or edge DP is too stringent to render useful performance guarantees for influence maximization. Similar observations have been made about the node and edge DP guarantees of other graph algorithms which are too sensitive to the presence of individual edges (and especially nodes) in their input graphs (Gupta et al. 2010, Li et al. 2022, Epasto et al. 2022, Blocki et al. 2013). In their analysis of deferentially private combinatorial optimization, Gupta et al. (2010) consider several graph problems including vertex and set cover. Their modified greedy algorithm for set cover is similar to our central DP mechanism in Section 3.1, although it addresses a substantially different privacy environment, where the goal is to protect a private subset $R$ while outputting a minimal collection that covers $R$. In both cases direct application of the exponential mechanism is inefficient and the greedy approach facilitates an efficient implementation.
@@
In formulating our notion of “influence sample differential privacy” (ISDP), we are moving away from having the social network as input to the algorithm. Instead, we focus on an input dataset of $m$ cascades that are collected at random on the social network graph. ISDP allows us to limit the information leakage about contribution of individual nodes to the input cascades (with local ISDP), or contribution of individual nodes to the algorithm output through the input cascades (with central ISDP). It should be noted that achieving local privacy under ISDP is easy by randomizing individual statuses of the nodes in the input cascades. However, in general, it is challenging to coordinate randomization between nodes in a meaningfully decentralized manner that matches the problem context, and the locally deferentially private analysis of graphs has been limited to certain restricted settings so far (Imola et al. 2021, Dhulipala et al. 2022).
@@
The fundamental relationship between information flow and privacy is well observed, cf. contextual integrity (Nissenbaum 2009, Benthall and Cummings 2024). However, the privacy implications of information diffusion and algorithmic intervention on social networks remain largely unexplored, with the exception of a few studies (Rezaei et al. 2021, Liell-Cock et al. 2020, Fanti et al. 2017). The existing work looks at this issue in specific, highly stylized scenarios. Rezaei et al. (2021) show that it is extremely difficult to hide the existence of a giant connected component of infected nodes under an independent cascade model, and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al. 2021). Liell-Cock et al. (2020) propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al. 2020). Motivated by the need to hide the identities of users who post sensitive messages, Fanti et al. (2017) propose a new messaging protocol called adaptive diffusion, which works by breaking the directional symmetry of diffusion around its true source and achieves fast spread and perfect source obfuscation for infinite regular trees, in the sense that all users with the message are equally likely to have been its originator. With our formulation of DP for influence maximization, we can focus on privacy leaks through the information flow in the network rather than the underlying network structure. Our formulation of DP for influence samples ($\epsilon$-ISDP in 1) allows us to explicitly quantify information leaks from the observation of random spreads on the network and to take a direct approach to the privacy risks that social network participants face from network diffusions.